# Post-detection inference for sequential changepoint localization

### v4

@@ 6 Experiments
Table 6.10: Comparison of the confidence sets for $T$ with [39] (Setting I)
$F_{0}$ $F_{1}$ $T$ Conditional Size Conditional Coverage Marginal Coverage
@@
The results show that [39] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable. However, our method (4.1) achieves the target conditional coverage.
@@
Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [31]. We want to localize the change from $60\%$ positive reviews to $40\%$ positive reviews. In our experimental setup, the true changepoint is at 500; the pre-change data are i.i.d. draws of $75\%$ positive reviews, and the post-change data are i.i.d. draws of $25\%$ negative reviews (these are all unknown to our algorithm). We construct a change detection method and a point estimate $\hat{T}$, using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [25], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations. After obtaining the binary sequence and $\hat{T}$, we construct the test statistic (3.4) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward). Note that given the change detector (and as a byproduct, the sequence of Bernoulli) and $\hat{T}$, it is simple to compute (3.4) in $O(T)$ time per $t$. So, the confidence set (3.9) takes total $O(T^{2})$ time. Fig. 6.11 provides a visualization.
@@
Figure 6.11: Vertical red dashed lines show the point estimate. The confidence set (3.9) is marked in red points. The true changepoint $T=500$ is marked with the black line. Texts are plotted from time $t=470$ upto detection time, $\tau=522$.
Remark 6.1.
@@
In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable. For nonparametric settings (IV, V, and VI), the only valid method currently known is the universal method. Prior works are valid (asymptotically) only for Setting I. As shown in Table 6.2 for known pre- and post-change distributions (Setting I), the adaptive method consistently outperforms the universal method, which tends to be conservative. Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplement H.1). However, for the composite settings (II and III), the universal method outperforms the adaptive method. Moreover, the universal method is also significantly more efficient — it runs in $O(T^{2})$ time (assuming $M_{t}$ can be computed in $O(T)$ time), while the adaptive method incurs a higher computational cost of $O(T^{3}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(T)$ time). See Supplement Sections E, F, and H for further experiments, involving confidence intervals for pre- and post-change parameters, detection procedures with bounded PFA, settings with dependent data, etc.
@@ 7 Conclusion
We developed a general framework for constructing confidence sets for the unknown changepoint $T$ after a sequential detector signals a change at a data-dependent stopping time. It applies to any general nonparametric post-change class when the pre-change distribution is known, and we further extend it to settings with a composite pre-change class under a suitable assumption. For the parametric setting, we proposed an alternative simulation-based method that empirically yields tighter confidence sets when both pre- and post-change distributions are known and can handle dependent data as well. However, we note that our primary (universal) method is nonparametric, computationally more efficient, and simpler to implement. Crucially, all our methods are agnostic to the detection algorithm used and provide non-asymptotic (conditional) coverage guarantees, whereas prior works are tailored for exponential family distributions with the CUSUM procedure and provide only asymptotic (conditional) coverage. By bridging a longstanding gap in the literature, our work offers the first nonparametric, theoretically principled, and broadly applicable method for post-detection inference.
@@
Outline of Supplementary Material
Omitted proofs can be found in Appendix A. We provide the implementation details of Algorithm 1 in Appendix B and of Algorithms 2 and 3 in Appendices C and D respectively, along with two concrete examples in both sections. Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix E. Experimental results for a different implementation method are in Appendix F. Simpler algorithms for PFA-controlling change detectors are presented in Appendix G. Some additional experiments with dependent (Markov) pre- and post-change data, with detection procedures with bounded PFA, and with Poisson rate change problems are in Appendix H. Some heuristic algorithms are presented in Appendices I and J, pointing at open avenues for further theoretical work.
@@ Appendix A Omitted proofs
Proof of Proposition 2.1..
For the sake of contradiction, assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
@@
Threfore, $\mathbb{P}_{\infty}(\tau=\infty)=1-\mathbb{P}_{\infty}(\cup_{n=1}^{\infty}[\tau<n])=1-\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau<n)=\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau\geq n)\geq 1-\alpha.$
Proof of Proposition 2.2.
@@
Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$. Since $\tau$ be a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$. This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ Therefore,
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C})$
$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C})$
$=1-\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
$\leq 1-(1-\alpha)\mathbb{P}_{F_{0},\infty}(\tau\geq t)$
@@
For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$
So, $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
@@
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T)$
@@
Lemma A.1 (Ville’s inequality [35]).
@@
For a short proof, see [11][Lemma 1].
Proof of Theorem 3.1.
@@
$=\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty]+\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty]$
$=\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{\hat{T}-1}\geq 2/\alpha,t<\hat{T}\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{\hat{T}}\geq 2/\alpha,\hat{T}<t\leq\tau<\infty\right]$
$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[S^{(t)}_{(t-1)\wedge\hat{T}}\geq 2/\alpha\right]$
Now, from Ville’s inequality (Lemma A.1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
@@
$\mathbb{P}_{F_{0},t,F_{1}}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha\right]$ $=\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{t\vee(\hat{T}-1)}\geq 2/\alpha,t\vee(\hat{T}-1)=n\right]\right)$
$\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
@@
Proof of Theorem 3.2.
@@
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<2/(\alpha r_{t})\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
@@
From Theorem 3.1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Now, since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\forall t\in\mathbb{N}.$ So, we obtain
@@
Proof of Theorem 3.3.
@@
Since $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}$ and $\hat{T}=\arg\max_{1\leq i\leq\tau}L_{i}^{\tau}$ and hence, whenever $\tau\geq T>t,$
$M_{t}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
Similarly, whenever $T<t\leq\tau,$ $M_{t}=\frac{L_{\hat{T}}^{\tau}}{L_{t}^{\tau}}\geq\frac{L_{{T}}^{\tau}}{L_{t}^{\tau}}=\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}$. Therefore,
@@
We also have
$|{\mathcal{C}}|\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
@@
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$ $\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
@@
$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$ $\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
@@
as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma A.2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely. Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{0}.$
@@
The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma A.3. Also, as we noted before, $\rho_{1}(s_{1})<1.$ Therefore,
@@
Proof of Proposition 3.7.
@@
where $X\sim P$ and $X^{\prime}\sim P^{*}$. By Theorem 1.A.2 of [26], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. Therefore,
@@
Proof of Theorem 3.8.
@@
The last steps follow from Assumption 3.5. The rest of the proof is the same as in the proof of Theorem 3.2. ∎
Proof of Theorem 3.9.
@@
Therefore, for any $t\leq T-1$, using the above inequality and the assumption (3.10a) we get
@@
In the last step, $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$ follows from (3.10a), $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (3.10c).
@@
First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$. Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$. So, for any fixed $c\in(0,1)$,
@@
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t)$ (S7)
@@
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality.
Proof of Theorem 4.4..
@@
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality.
Proof of Theorem 5.1.

### v6

@@ 6 Experiments
Table 7: Comparison of the confidence sets for $T$ with [wu2007inference] (Setting I)
$F_{0}$ $F_{1}$ $T$ Conditional Coverage Conditional Size Marginal Coverage
@@
The results show that [wu2007inference] (which is only asymptotically valid and relies on approximations) exhibits noticeable undercoverage, while the sizes for both methods are comparable. However, our method (17) achieves the target conditional coverage.
@@
Sentiment change in language data: We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment [socher2013recursive]. We want to localize the change from $60\%$ positive reviews to $40\%$ positive reviews. In our experimental setup, the true changepoint is at 500; the pre-change data are i.i.d. draws of $75\%$ positive reviews, and the post-change data are i.i.d. draws of $25\%$ negative reviews (these are all unknown to our algorithm). We construct a change detection method using a pre-trained DistilBERT-base model fine-tuned on the uncased SST-2 [sanh2019distilbert], which is a binary sentiment classifier that maps each review to either 0 (negative) or $1$ (positive), thereby producing a sequence of Bernoulli observations. After obtaining the binary sequence, we construct the test statistic (9) using the likelihood ratio process for $\text{Bern}(0.4)$ against $\text{Bern}(0.6)$ (forward) and for $\text{Bern}(0.6)$ against $\text{Bern}(0.4)$ (backward). Note that given the change detector (and as a byproduct, the sequence of Bernoulli), it is simple to compute (9) in $O(T)$ time per $t$. So, the confidence set (14) takes total $O(T^{2})$ time. Fig. 4 provides a visualization.
@@
Figure 4: The confidence set (14) is marked in red points. The true changepoint $T=500$ is marked with the black line. Texts are plotted from time $t=470$ upto detection time, $\tau=534$.
Remark 4.
@@
In Settings I, II, and III, both of our proposed methods — universal (Section 3) and adaptive (Section 4) — are applicable. For nonparametric settings (IV and V), the only valid method currently known is the universal method. Prior works are valid (asymptotically) only for Setting I. As shown in Table 2 for known pre-distributions (Settings I and II), the adaptive method consistently outperforms the universal method, which tends to be conservative. Also, our adaptive method is the only method we know that remains valid under dependence (see experimental results in Supplementary Section G.2). However, for setting III, the universal and adaptive methods perform similarly. Moreover, the universal method is also significantly more efficient — it runs in $O(T^{2})$ time (assuming $M_{t}$ can be computed in $O(T)$ time), while the adaptive method incurs a higher computational cost of $O(T^{3}B)$ (assuming $\mathcal{A}$ runs in $O(T)$ time and $M_{t}$ can be computed in $O(T)$ time). See Supplement Sections E, F, and G for further experiments, involving confidence intervals for pre- and post-change parameters, settings with dependent data, settings with detection algorithms controlling PFA, etc.
@@ 7 Conclusion
We developed a general framework for constructing confidence sets for the unknown changepoint $T$ after a sequential detector signals a change at a data-dependent stopping time. It applies to any general nonparametric post-change class when the pre-change distribution is known, and we further extend it to settings with a composite pre-change class under a suitable assumption. For the parametric setting, we proposed an alternative simulation-based method that empirically yields tighter confidence sets when both pre- and post-change distributions are known and can handle dependent data as well. However, we note that our primary (universal) method is nonparametric, computationally more efficient, and simpler to implement. Crucially, all our methods are agnostic to the detection algorithm used and provide non-asymptotic (conditional) coverage guarantees, whereas prior works are tailored for exponential family distributions with the CUSUM procedure and provide only asymptotic (conditional) coverage. By bridging a longstanding gap in the literature, our work offers the first nonparametric, theoretically principled, and broadly applicable method for post-detection inference. A particularly fruitful direction for future work is to extend this framework to the distribution-free setting.
Acknowledgement
Part of this work was carried out while AS was a student at the Indian Statistical Institute. The authors thank the Editor, the Associate Editor, and the reviewers for their careful reading of an earlier version of the article and for their helpful comments. The authors also thank Swapnaneel Bhattacharyya for identifying and helping eliminate an inconsistency in an earlier version of the article. AR acknowledges support from NSF DMS-1916320.
@@ Acknowledgement
Outline of the Supplementary Material
Omitted proofs can be found in Appendix A. We provide the implementation details of Algorithm 1 in Appendix B and of Algorithms 2 and 3 in Appendices C and D respectively, along with two concrete examples in both sections. Experimental results for the confidence intervals for pre- and post-change parameters are in Appendix E. Simpler algorithms for PFA-controlling change detectors are presented in Appendix F, along with their implementations in simulation experiments. Additional experiments with varying $T$ vs. ARL and scenarios with dependent (Markov) pre- and post-change data are presented in Appendix G.
@@ Appendix A Omitted proofs
Proof of Proposition 1..
Assume that $\mathbb{P}_{T}(T\in\mathcal{C}^{*})\geq 1-\alpha$ for some $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$. Now we observe that
@@
Therefore, $\mathbb{P}_{\infty}(\tau=\infty)=1-\mathbb{P}_{\infty}(\cup_{n=1}^{\infty}[\tau<n])=1-\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau<n)=\lim_{n\to\infty}\mathbb{P}_{\infty}(\tau\geq n)\geq 1-\alpha.$
Proof of Proposition 2..
@@
Let $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying $\mathbb{P}_{F_{0},t,F_{1}}(t\in\mathcal{C}\mid\tau\geq t)\geq 1-\alpha$, for all $t\in\mathbb{N}$. Since $\tau$ is a stopping time, $\tau<t]=\displaystyle\cup_{i=1}^{t-1}[\tau=i]$ does not depend on $\{X_{i}:i\geq t\}$. This implies $\mathbb{P}_{F_{0},\infty}(\tau\geq t)=\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t).$ Therefore,
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)$ $=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C},t\leq\tau)$
$=\mathbb{P}_{F_{0},t,F_{1}}(t\notin\mathcal{C}\mid\tau\geq t)\mathbb{P}_{F_{0},t,F_{1}}(\tau\geq t)$
@@
For the converse part, $\phi^{(t)}$ satisfies $\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$. So, $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
@@
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})+\mathbb{P}_{F_{0},\infty}(\tau<T).$
@@
This completes the proof. ∎
Lemma 1 (Ville’s inequality [ville1939etude]).
@@
For a short proof, see [howard2020time][Lemma 1].
Proof of Theorem 1..
@@
$=\mathbb{P}_{F_{0},t,F_{1}}\left[\max\{\max_{t\leq n\leq\tau}R^{(t)}_{n},\max_{1\leq n\leq t-1}S^{(t)}_{n}\}\geq 2/\alpha,t\leq\tau<\infty\right]$
$\leq\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{t\leq n\leq\tau}R^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right]+\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{1\leq n\leq t-1}S^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right].$
Now, from Ville’s inequality (Lemma 1) on the e-processes $\{R^{(t)}_{n}\}_{n\geq t}$,
@@
$\mathbb{P}_{F_{0},t,F_{1}}\left[\max_{t\leq n\leq\tau}R^{(t)}_{n}\geq 2/\alpha,t\leq\tau<\infty\right]$ $\leq\mathbb{P}_{F_{0},t,F_{1}}\left(\cup_{n=t}^{\infty}\left[R^{(t)}_{n}\geq 2/\alpha\right]\right)$
@@
This completes the proof. ∎
Proof of Theorem 2..
@@
$=\mathbb{P}_{F_{0},T,F_{1}}(T\in\{t\in\mathbb{N}:t\leq\tau,M_{t}<\frac{2}{\alpha r_{t}}\}\mid\tau\geq T)\mathbb{P}_{F_{0},T,F_{1}}(\tau\geq T)+\mathbb{\mathbb{P}}_{F_{0},T,F_{1}}(\tau<T)$
@@
From Theorem 1, we have $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]\leq\alpha r_{t}$, since $r_{t}$ and $M_{t}$ are independent of each other. Now, since $r_{t}$ is an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}]=\mathbb{E}[\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq\frac{2}{\alpha r_{t}}\mid r_{t}]]\leq\mathbb{E}[\alpha r_{t}]=\alpha\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, $\forall t\in\mathbb{N}.$ So, we obtain
@@
This completes the proof. ∎
Proof of Theorem 3..
@@
Whenever $\tau\geq T>t,$
$M_{t}\geq\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}.$
Similarly, whenever $T<t\leq\tau,$
$M_{t}\geq\prod_{i=T}^{{t}-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$
Therefore, the following holds almost surely:
@@
Also, the following holds almost surely:
$|{\mathcal{C}}|\times\mathds{1}(\tau\geq T)\leq\sum_{t=1}^{{T-1}}\mathds{1}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)+1+|\tau-T|.$ (S2)
@@
$\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
$\leq\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{T})\right)\quad\text{[Since }r_{t}\leq r_{T}]$
@@
$\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\mid\tau\geq T\right)$
$\leq\frac{1}{p_{T}}\sum_{t=1}^{{T-1}}\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=t}^{{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\leq 2/(\alpha r_{t})\right)$
@@
as $N\to\infty.$ To justify the convergence, we note that $\left(\frac{2}{\alpha r_{T}}\right)^{s}$ is uniformly integrable as a sequence in $N\in\mathbb{N}$ (follows from Lemma 2), and by the strong law of large numbers, $\left(\frac{2}{\alpha r_{T}}\right)^{s}\to\left(\frac{2}{\alpha p_{T}}\right)^{s}$ almost surely. Since the above holds for any $s\in(0,1)$, it also holds for $s=s_{0}.$
@@
The last step follows from the assumption that $\tau$ is sensible (recall that if $\tau$ is sensitive, then $\tau_{j}$ is stochastically larger than $\tau$) and from Lemma 3. Also, as we noted before, $\rho_{1}(s_{1})<1.$ Therefore,
@@
This completes the proof. ∎
Proof of Proposition 3..
@@
where $X\sim P$ and $X^{\prime}\sim P^{*}$. By Theorem 1.A.2 of [shaked2007stochastic], the stochastic ordering implies that there exist a random variable $Z$ and functions $\psi_{1}$ and $\psi_{2}$ such that $\psi_{1}(z)\leq\psi_{2}(z)$ for all $z$ and $\frac{p_{1}^{*}(X)}{p_{0}^{*}(X)}\stackrel{{\scriptstyle d}}{{=}}\psi_{1}(Z)$ and $\frac{p_{1}^{*}(X^{\prime})}{p_{0}^{*}(X^{\prime})}\stackrel{{\scriptstyle d}}{{=}}\psi_{2}(Z)$. Therefore,
@@
This completes the proof. ∎
Proof of Theorem 4..
@@
The last steps follow from Assumption 1. The rest of the proof is the same as in the proof of Theorem 2. ∎
Proof of Theorem 5..
@@
Therefore, for any $t\leq T-1$, using the above inequality and the assumption (15a) we get
@@
In the last step above, we used the inequality $\mathbb{P}_{F_{0},\infty}(\tau\geq T)\geq\mathbb{P}_{F_{0}^{*},\infty}(\tau\geq T)$, which follows from (15a), and $\mathbb{P}_{F_{0},T,F_{1}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ $\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}\left(\prod_{i=T}^{{t}-1}\frac{f_{1}^{*}(X_{i})}{f_{0}^{*}(X_{i})}\leq 2/(\alpha\eta)\right)$ follows from the fact that $(F_{0}^{*},F_{1}^{*})$ is an LFD pair, and $\mathbb{P}_{F_{0},T,F_{1}}(r_{\tau}\leq\eta)\leq\mathbb{P}_{F_{0}^{*},T,F_{1}^{*}}(r_{\tau}\leq\eta)$ follows from (15c).
@@
First, observe that $M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}$ are i.i.d. under $H_{0,t}$. Hence, $M_{t}\leq\operatorname{Quantile}(1-c;M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B})$ $\iff M_{t}$ is one among the $\lceil(1-c)(B+1)\rceil$ smallest of the collection $\{M_{t},M_{t,\infty}^{1},\cdots,M_{t,\infty}^{B}\}$. So, for any fixed $c\in(0,1)$,
@@
$\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,L})\leq\mathbb{P}_{F_{0},t,F_{1}}(M_{t}>C_{t,\infty})\leq\alpha\mathbb{P}_{F_{0},\infty}(\tau\geq t).$ (S7)
@@
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{T,L}^{\alpha r_{t}})\geq 1-(\alpha+\beta)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta$ from the above inequality. ∎
Proof of Theorem 8..
@@
Since $\mathbb{P}_{\theta_{0},T,\theta_{1}}(M_{T}\leq C_{t,L})\geq 1-(\alpha+\beta+\gamma)\times\mathbb{P}_{\theta_{0},\infty}(\tau\geq T)$, we obtain $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma$ from the above inequality. ∎
Proof of Theorem 9..